Please help with a detailed solution.

Problem 2 (1). (35 pts total) 2-0 The inlet volumetric flow rate for an adiabatic compressor is 10 (L/s] of air. The inlet co(a) (5 pts) What is the heat transfer into the system? v. 2.001m3 Answer = Answer = _ 0 _[kJ/kg] Valls (b) (5 pts) What is
Answer #1

Given that in let volumetric flow rate (g) = 10 Hs 9 = 10 45 = 10x 103 m/s (11-103 m) 9 = 0.01 m3/s. Inlet conditions P = 12(d) Mass flow rate (n) = 9 2, 0.01 0.7 Ñ = 0.014285 kgfs (es Increase in specific enthalpy of air from inlet to outlet is th